OpenStudy (anonymous):

log4(x-4)+log4(x+59)=3

OpenStudy (solomonzelman):

\(\Large\color{black}{ \displaystyle \log_\color{red}{\rm a}\left( \color{blue}{\rm b} \right)+\log_\color{red}{\rm a}\left( \color{blue}{\rm d} \right)=\log_\color{red}{\rm a}\left( \color{blue}{\rm b \times d } \right) }\)

OpenStudy (solomonzelman):

use that for your left side.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \log_\color{red}{\rm 4}\left( \color{blue}{\rm x-4} \right)+\log_\color{red}{\rm 4}\left( \color{blue}{\rm x+59} \right)=\log_\color{red}{\rm a}\left( \color{blue}{\rm ~~~(x-4)\times (x+59)~~~} \right) }\)

OpenStudy (solomonzelman):

the last "a" on the right should be 4

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \log_\color{red}{\rm 4}\left( \color{blue}{\rm ~~~(x-4)\times (x+59)~~~} \right)=3 }\) \(\large\color{black}{ \displaystyle \log_\color{red}{\rm 4}\left( \color{blue}{\rm ~~x^2-4x+59x-236~~} \right)=3 }\) \(\large\color{black}{ \displaystyle \log_\color{red}{\rm 4}\left( \color{blue}{\rm ~~x^2+55x-236~~} \right)=3 }\) \(\large\color{black}{ \displaystyle x^2+55x-236=4^3 }\) \(\large\color{black}{ \displaystyle x^2+55x-236=64 }\) \(\large\color{black}{ \displaystyle x^2+55x-300=0 }\)

OpenStudy (anonymous):

would i do the quadratic formula next?

OpenStudy (solomonzelman):

factor it

OpenStudy (solomonzelman):

60 - 5 = 55 60 * (-5) = -300

OpenStudy (anonymous):

(x+60)(x-5)= -60,5